OpenStudy (jenniferjuice):

If f(x) = -2x + 1, find f(-6)

OpenStudy (jenniferjuice):

12 13 -11 -7

OpenStudy (anonymous):

f(x) = -2x + 1 f(x) ==> f(-6) Substitute (-6) for "x" values f(-6) = -2(-6) + 1 f(-6) = 12 + 1 f(-6) = ___

OpenStudy (anonymous):

@jenniferjuice do you understand? what did you get?

OpenStudy (jenniferjuice):

um well not really ...

OpenStudy (jenniferjuice):

@some_someone

OpenStudy (anonymous):

Okay, what are you uncertain on? Any specific step(s)?

OpenStudy (marissalovescats):

You just plug in -6 for the x :)

OpenStudy (jenniferjuice):

what do i do at the end? i got the substituting part

OpenStudy (anonymous):

^ yup I said that and demonstrated that :) All he had to do at the end is just simply do some adding and get the answer, since i did step by step.

OpenStudy (jenniferjuice):

f(-6) = ___

OpenStudy (anonymous):

f(-6) = -2(-6) + 1 f(-6) = 12 + 1 f(-6) = ___ *Just add*

OpenStudy (jenniferjuice):

oh lol i thought i was supposed to divide ..

OpenStudy (anonymous):

I left a blank so that you could put the answer.

OpenStudy (jenniferjuice):

okay so f(-6)=13

OpenStudy (anonymous):

yup :)

OpenStudy (jenniferjuice):

so B?

OpenStudy (anonymous):

B.) 13 Seems right.

OpenStudy (jenniferjuice):

okay thankyouuu ! :)

OpenStudy (anonymous):

yw :)